Q1. 100 ml of water sample on titration with N/50 HCl requires 8 mL of acid for phenolphthalein end-point and 9 mL of total acid for methyl-orange end-point. Calculate the type and extent of alkalinity present in water sample.
1
Expert's answer
2019-10-03T05:19:17-0400

Here diacidic base is present which give two end point one with phenalphthein and second with methy orange

For phenolphthalein end point-

"\frac{1}{2}\ Meq\ of\ Diacidic\ base = Meq.\ of\ of \ 8ml \ \frac{N}{50}\ HCl"

="\frac{1\times 8}{50\times1000}=\frac{1}{6225}\ Meq\\"

"Meq.\ of\ Diacidic\ base=2\times Meq.\ of\ HCl=\frac{2}{6225}\ Meq."

For methyl orange end point-

"Meq.\ of\ monoacidic\ base\ formed=Meq.\ of\ 1ml of\ extra\ HCl\ used"

"=\frac{1\times1}{50\times 1000}=\frac{1}{50000\ Meq.}"


Normality of diacidic alkalinity present in 100 ml water = "\frac{2\times 1000}{6225\times 100}=\frac{20}{6225}=\frac{4}{1245}\ N"